OpenStudy (anonymous):

Which regular polygon would have each of its interior angles measure 120°? hexagon heptagon octagon nonagon

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large{180(n - 2) \over n} = 120 }\) With this way, you'd determine the number of sides(that is \(n\))

OpenStudy (anonymous):

180(n-2)=n 120 n=6

OpenStudy (anonymous):

120+360/n=180 360/n=180-120=60 60*n=360 n=360/60=6 hexagon

OpenStudy (unklerhaukus):